complete combustion of 8.00g a hydrocarbon produced 26.0g of CO2 and 7.99g of H2O. what is the empirical formula for the hydrocarbon?
1
Expert's answer
2019-09-30T05:15:33-0400

Consider the general balanced equation for combustion of hydrocarbon:

"C_xH_y+(x+\frac{y}{4})O_2\to xCO_2+\frac{y}{2}H_2O"

Molar mass of "CO_2=44\frac{g}{mol}"

So the number of moles of "CO_2" formed"=x=\frac{26}{44}=0.59""mol"

Molar mass of "H_2O=18\frac{g}{mol}"

So the number of moles of "H_2O" formed"=\frac{y}{2}=\frac{7.99}{18}=0.44""mol"

So "y=2\times0.44=0.88mol"

For "C: x=0.59\approx0.60"

For"H:y=0.88\approx0.90"

So there will be 3 "H" atoms per 2 "C" atoms

Hence the empirical formula of the given hydrocarbon is "C_2H_3"